What number comes next and why? 2,3,6,15,42,_____
Answer by scott8148(6628)

123

look at the sequence of the differences between consecutive terms

3 - 2 = 1 = 3^0
6 - 3 = 3 = 3^1
15 - 6 = 9 = 3^2
42 - 15 = 27 = 3^3

123 - 42 = 81 = 3^4